# David and Goliath: An Empirical Evaluation of Attacks and Defenses for QNNs at the Deep Edge

### v1

@@ 1 Introduction
The availability of large volumes of data and tremendous computing power turned machine learning (ML) into a critical technology to drive innovation and assist humans in tasks ranging from healthcare [1] and finance [2] to mobility [3]. As ML workloads tend to be computationally intensive, most ML systems developed to date followed a cloud-centered design, sending the data collected at the edge to the cloud [4, 5, 6]. The cloud performs the ML computation and returns the final decision to the edge responsible for actuating over the environment. However, recent concerns about privacy, latency, power consumption, and cost are forcing ML to move from the cloud to the edge near the data source [7, 8].
Shifting ML workloads to the edge reduces the exposure of private data, as there is no data transmission over the network [7, 8]. Besides reducing the chances of data stealing, this also increases the users’ confidence in applications and services as it lessens the need to provide personal data to a third party. Furthermore, edge computing enables ML systems to be more predictable as the latency is not dependent on the available network bandwidth [6, 7, 8]. Therefore, edge computing enables developers to provide reliable throughput guarantees, which are mandatory in real-time or mission-critical scenarios [9].
Edge devices are now the new frontier for ML deployment [6, 7, 8, 10]. Of the panoply of devices typically deployed at the edge of the network, microcontrollers (MCUs) are getting special attention, as newer architectures aligned with friendly developer tools (e.g., CMSIS-NN [11], PULP-NN [12], and TensorFlow (TF) Lite Micro [13]) are enabling the deployment of ML systems in devices that are several orders of magnitude cheaper and less power-hungry than application processors (APUs) or graphical processing units (GPUs). Despite the resource-constrained nature of these devices, artificial neural network (ANN) compression techniques such as quantization have proved to have negligible impact on accuracy while tremendously improving latency [11, 12].
In parallel with this development, recent research has shown that full-precision ANNs are vulnerable to the so-called adversarial examples [14, 15, 16], i.e., inputs carefully crafted to fool the neural network into making an obvious misclassification. In this context, it is essential to ask if quantized neural networks (QNNs) are susceptible to this class of attacks and, if so, how QNNs could be protected. Previous research found that QNNs are still vulnerable to gradient-free attacks and adversarial example transferability, even to those crafted on full-precision ANNs using weak attacks such as Fast Gradient Sign Method (FGSM) [7, 8, 17, 18, 19]. [18] pioneered in adversarial defenses for QNNs by proposing a novel training algorithm that reduces the quantization-shift phenomenon; however, the authors do not provide a clear picture of the scope of their work. As GPUs operate better with integer precision, QNNs are deployed not only on ultra-low-power MCUs but also on powerful GPUs. As GPUs have far more memory and computing power than ultra-low-power MCUs, QNNs deployed on these devices show significant differences in the quantization policy, kernel filter size, and network depth. In this context, providing a clear picture of the quantization policy and the target computing devices when working with QNNs is fundamental. To the best of our knowledge, no work proposes possible defense mechanisms or evaluates the feasibility of adapting the existing defenses for full-precision ANNs on QNNs tailored for the TinyML scenario.
@@ 2.1 TinyML: The New Wave in AI
TinyML is a field of ML dedicated to the design and development of hardware and software solutions to enable the execution of ML inference under a milliWatt scale [20]. By enabling on-device and near-sensor data analysis, TinyML eliminates the imperative reliance on the cloud infrastructure for inference results [7, 8]. This paradigm shift not only enables the dissemination of ML in regions without reliable network, but also ensures greater responsiveness and bolstered privacy.
Under this umbrella, ML at the extreme edge (powered by MCUs) is one of the hottest topics in the ML field due to the proliferation of these devices in our infrastructures/society. In effect, ARM predicts that MCUs are expected to reach 1 trillion across various market segments by 2035 [11]. A recent report [21] from ABI Research dating to 2022 emphasized that TinyML Software-as-a-Service was expected to exceed US\$220 million in revenue in 2022 and have the potential to become a billion-dollar market by 2030. This same company also forecasts that the TinyML market will grow from 15.2 million shipments in 2020 to 2.5 billion in 2030 [22].
In academia, this trend has fueled the development of multiple software and hardware solutions to enable and optimize the execution of TinyML and has fuelled the rise of RISC-V instruction set architecture (ISA). Among other relevant works, [23] proposed a custom SoC based on the emerging RISC-V ISA [23], while [12] proposed a software library that optimizes neural network functions on this same SoC. [24] proposed a RISC-V hardware accelerator able to deliver 3200 inferences/second while only consuming 12.2 milliWatts.
In terms of use cases, TinyML has been widely deployed for (i) word spotting, (ii) object recognition, (iii) object counting, and (iv) audio/voice detection [20, 21]. More recently, its application has been extended to the detection of (i) forest fire, (ii) shape, and (iii) seizure [21].
@@ 2.2 Scope
This work targets ultra-low-power ML applications (TinyML), commonly powered by Arm Cortex-M (Armv7-M or Armv8-M) and RISC-V MCUs (e.g., PULP [25]). These devices typically have a single core and feature Flash and static random access memories (SRAM). They execute instructions in place from physical memory and lack a memory management unit (MMU) for virtual memory support. Consequently, they are unable to run Unix-like operating systems.
@@
To help standardize the throughput of ultra-low-power hardware on ML applications, [20] developed the MLPerf Tiny benchmark, composed of four QNNs targeting image and sound applications. Our empirical evaluation targets QNNs developed under the exact specification as MLPerf Tiny: QNNs using fixed-point quantization with 8 or 16 bits for weights and activations and quantized according to the TF Lite Micro policy. Within this policy, the number of bits representing the integer and fractional parts is not fixed, varying for weights and activations within the same layer.
@@ 2.3 Adversary Model
Adversary Goal. The adversary’s goal is to compromise the decision integrity of a supervised model, disrupting severe accuracy drops. Previous research shows that the attacker can achieve this goal by interfering with the training or inference phase. However, research on training at the deep edge is still at an embryonary stage [4], and this class of devices is still only used for inference purposes. The most relevant part of the training of the models deployed on deep-edge devices is still performed in cloud-like environments. Therefore, we limit the scope of our research to the inference phase. Attacks to craft adversarial examples can either be targeted or non-targeted. Targeted attacks misguide an ANN by making it classify an adversarial sample as a specific class. In contrast, in non-targeted attacks, the class returned by the model is arbitrary except for the original one. Our research encompasses only non-targeted attacks, as they enable a broader assessment of the robustness and generalization of a neural network.
Adversary Capability. Capability refers to the attacker’s capacity to control or manipulate the input data. At inference, the generic data processing pipeline of an ML system considers that the (i) input features are collected from sensors or data repositories and (ii) processed in the digital domain, (iii) used by the model to infer about the surrounding environment, (iv) which communicates the result to the appropriate actuator [9]. Considering this pipeline, adversarial examples can be generated either in the physical domain, by tampering with sensors, or in the digital domain. This work only considers attacks and defenses in the digital domain, which dominates the available literature.
@@ 2.4 Selected Datasets and Models
To benchmark the throughput of the hardware used in TinyML, [20] proposed the MLPerf Tiny benchmark. This benchmark results from the collaborative effort of more than 30 organizations from industry and academia and reflects the community’s needs by simulating real-world applications (see Section 2.1). As the datasets and models used in these applications are the companies’ intellectual property, researchers selected representative and open-sourced models and datasets.
@@
We included a third dataset because evaluating attacks and defenses on only two datasets may fall short. For this purpose, we borrowed the work [26], which lists and reviews ML implementations for Arm Cortex-M MCUs. After analyzing the reproducibility of each listed work, we replicated the work [27]. This work addressed the development of an ML system devoted to detecting coffee plant diseases. The dataset is public and contains 1747 images for training and testing distributed over four classes, each corresponding to a different disease. All QNNs included in our study are quantized post-training.
@@ 2.5 Selected Attacks and Defenses
We prioritize attacks/defenses with great impact and citation in the research community and whose implementation is open-sourced. Furthermore, since we target QNNs following the TF Lite Micro quantization policy, we prioritize attacks and defenses implemented in TensorFlow 2 (TF2) or Keras (the backend engine of TF2). TF2 provides straightforward mechanisms to operate with models in TF Lite format. We also prioritize attacks/defenses available in ART [28], Cleverhans [29], or Foolbox [30, 31] frameworks, as these frameworks provide a straightforward mechanism to benchmark different attacks and defenses upon a common interface. Our evaluation suite includes six white-box, two gray-box, and two black-box attacks, described in Table II.
@@
Defenses. We consider six defense mechanisms (Table III) distributed across the two types of defense proposed by the research community. Our empirical evaluation includes four train-based defenses and the other two based on input pre-processing. Although there might be more recent defenses available in the literature than those we selected, PGD Adversarial Training, proposed in 2018 and included in our work, remains the state-of-the-art defense against adversarial attacks according to recent studies [32, 33, 34]. Nevertheless, we still include a recent iteration of this defense, entitled Sinkhorn Adversarial Training (2021), which promises to reduce the ground distance between original images and their adversarial representations.
@@
[…]
@@
Adversarial training consists in retraining an ANN by injecting adversarial examples into the training dataset. The first work on adversarial training [45] proposed to generate adversarial examples using FGSM in each training iteration, mixing clean samples and adversarial examples. This work was later borrowed by [38], which proposed adversarial training with PGD.
@@ 2.7 Threats to Validity
The experiments were conducted by two independent researchers on two separate but equal platforms: STM NUCLEO-L552ZE-Q. Whenever possible, we used the implementation of attacks and defenses available on the ART framework [28], except for AutoAttack and Sinkhorn Adversarial Training. Regarding AutoAttack, the original implementation is significantly different from the implementation of ART. Sinkhorn Adversarial Training is not available on ART and is originally implemented in PyTorch. In these two cases, we borrowed the original implementations and adapted them to work with TF2. Despite all efforts, there may be slight variations compared to the original works. To encourage further reproducibility of our results, all artifacts are available open-source [^1] https://gitlab.com/ESRGv3/QNN-Adversarial-Robustness.git.
@@ 3.1 Direct Attacks
The fact that the distortion significantly increases for similar or higher adversarial accuracy as the precision lowers suggests that the average point distance to the decision boundary increases as precision lowers. This finding is reinforced in the analysis A5 of Section 3.2 and was previously observed in [8]. [8] defined a geometric model proving that the average distance to the decision boundary increases as the weight precision and the input dimensionality decrease.
Takeaway 1. The average point distance to the decision boundary increases as the precision of a neural network decreases, making int-8 models generally more robust to adversarial examples than int-16 and float-32 versions.
@@
[7] arrived at a similar conclusion. They observed that the quantization of activations could make the gradient explode or vanish. However, they did not observe the same behavior for weight-only quantized models, i.e., models with weights quantized to integer precision but whose activations remain in floating-point.
Takeaway 2. Quantization increases the chance of gradient masking. Small perturbations may be erased by quantization and make the gradient vanish while bigger perturbations can make the gradient explode.
@@ 3.2 Transferability
A3. Transferability is sensitive to quantization-shift. As observed in Figure 1, for the CIFAR-10 and Coffee datasets, the lines representing the adversarial accuracy of float-32 and int-16 models overlap for the entire range of distortion. However, the same does not occur with the int-8 model. While for float-32 and int-16, the adversarial accuracy decreases smoothly as distortion increases, for the int-8 model, the adversarial accuracy decreases in buckets - it only decreases after a certain distortion threshold. We argue that this is a consequence of the quantization shift, the most widely-known effect of quantization [7, 8, 17, 18, 19, 50]. As quantization maps more than one value on the same quantization bucket, quantization can amplify or attenuate the effect of distortion. Quantization can boost adversarial accuracy by eliminating small perturbations; however, it can amplify bigger perturbations.
@@
Takeaway 3. Adversarial example transferability is influenced by quantization-shift. Quantization can mitigate the impact of small perturbations while amplifying the effects of larger perturbations.
A4. Gradient misalignment. In addition to the quantization-shift phenomenon, the poor transferability results generally observed from float-32 to int-16 and int-8 models can also be a consequence of gradient misalignment, as previously disclosed by [7]. Quantization can map multiple float-32 values into the same int-16 or int-8 value, moving distinct ANN weights and activations into the same quantization bucket. Therefore, the gradient estimated for float-32 ANNs may differ for the equivalent int-16 or int-8 models and even have different directions, with some attacks being more affected than others by this phenomenon.
@@
Takeaway 4. By mapping different ANN weights and activations into the same integer value, quantization can cause gradient misalignment: the gradient estimated for float-32 ANNs may not be the same for the equivalent int-16 or int-8 models and even have different directions.
@@ 4 Defenses Against Adversarial Examples
  • Gray-box: We only include the Square Attack as the results detailed in Tables V suggest that ZOO does not perform well against the neural networks we are using.
@@ 4.1 Train-Based Defenses
Takeaway 5. PGD and Sinkhorn Adversarial Training consistently increase the average point distance to the decision boundary, and this feature holds after quantization. Defensive Distillation and Ensemble Adversarial Training are not consistent in this topic.
@@
Takeaway 6. None of the training-based defenses mitigate adversarial example transferability. We envision this is a consequence of none of them smoothing the quantization-shift and gradient misalignment, which are required to tackle the transferability.
@@
Takeaway 7. Every time a defense is effective in a full-precision ANN, it is likewise effective in the corresponding QNNs. A defense may, however, be adequate to defend a low-precision QNN but insufficient to defend a full-precision ANN.
@@ 4.2 Pre-Processing Defenses
Pre-processing defenses are evaluated against the adversarial examples produced to attack the vanilla ANNs and QNNs before any defense. We do not craft new adversarial examples for the Defense+ANN or Defense+QNN system as pre-processing defenses are known to disrupt the well-known gradient-masking phenomenon, in which the gradient becomes useless to create an adversarial example, giving a false sense of security [7, 8, 18]. The evaluation of Pixel Defend was restricted to CIFAR-10 due to hardware limitations: PixelCNN is a heavy network whose amount of neurons grows proportionally with the size of input images. Therefore, we limited the evaluation scope to the dataset with smaller images. The configuration of Feature Squeezing and Pixel Defend is detailed in Table XV (Appendix A), while the test accuracy of the ANNs/QNNs enhanced with these defenses is detailed in Table X. Tables XI and XII detail the results observed for Feature Squeezing and Pixel Defend, respectively.
@@
Takeaway 8. Feature Squeezing can denoise a larger fraction of adversarial examples than Pixel Defend. In contrast to Pixel Defend, Feature Squeezing is portable to edge devices powered by Arm Cortex-M MCUs as long as they feature a DSP.
@@
Takeaway 9. Feature Squeezing is effective against small perturbations, but the spatial smoothing filter starts to fall short as the amount of perturbed pixels increases.
@@
Takeaway 10. Train-based defenses do not incur additional computation during inference and, therefore, are preferable over pre-processing defenses to be deployed in edge devices powered by Arm Cortex-M MCUs.

### v2

@@ 1 Introduction
The availability of large volumes of data and tremendous computing power turned machine learning (ML) into a critical technology to drive innovation and assist humans in tasks ranging from healthcare [miotto2017] and finance [ahmed2022] to mobility [ying2021]. As ML workloads tend to be computationally intensive, most ML systems developed to date followed a cloud-centered design, sending the data collected at the edge to the cloud [costa2022, murshed2021, stoica2017]. The cloud performs the ML computation and returns the final decision to the edge responsible for actuating over the environment. However, recent concerns about privacy, latency, power consumption, and cost are forcing ML to move from the cloud to the edge near the data source [bernhard2019, gorsline2021].
Shifting ML workloads to the edge reduces the exposure of private data, as there is no data transmission over the network [bernhard2019, gorsline2021]. Besides reducing the chances of data stealing, this also increases the users’ confidence in applications and services as it lessens the need to provide personal data to a third party. Furthermore, edge computing enables ML systems to be more predictable as the latency is not dependent on the available network bandwidth [stoica2017, bernhard2019, gorsline2021]. Therefore, edge computing enables developers to provide reliable throughput guarantees, which are mandatory in real-time or mission-critical scenarios [papernot2018].
Edge devices are now the new frontier for ML deployment [stoica2017, bernhard2019, gorsline2021, costa2022_2]. Of the panoply of devices typically deployed at the edge of the network, microcontrollers (MCUs) are getting special attention, as newer architectures aligned with friendly developer tools (e.g., CMSIS-NN [cmsis-nn], PULP-NN [pulp-nn], and TensorFlow (TF) Lite Micro [tflitemicro]) are enabling the deployment of ML systems in devices that are several orders of magnitude cheaper and less power-hungry than application processors (APUs) or graphical processing units (GPUs). Despite the resource-constrained nature of these devices, artificial neural network (ANN) compression techniques such as quantization have proved to have negligible impact on accuracy while tremendously improving latency [cmsis-nn, pulp-nn].
[…]
@@ 2.1 TinyML: The New Wave in AI
TinyML is a field of ML dedicated to the design and development of hardware and software solutions to enable the execution of ML inference under a milliWatt scale [tinyml]. By enabling on-device and near-sensor data analysis, TinyML eliminates the imperative reliance on the cloud infrastructure for inference results [bernhard2019, gorsline2021]. This paradigm shift not only enables the dissemination of ML in regions without reliable network, but also ensures greater responsiveness and bolstered privacy.
Under this umbrella, ML at the extreme edge (powered by MCUs) is one of the hottest topics in the ML field due to the proliferation of these devices in our infrastructures/society. In effect, ARM predicts that MCUs are expected to reach 1 trillion across various market segments by 2035 [cmsis-nn]. A recent report [abi2022] from ABI Research dating to 2022 emphasized that TinyML Software-as-a-Service was expected to exceed US\$220 million in revenue in 2022 and have the potential to become a billion-dollar market by 2030. This same company also forecasts that the TinyML market will grow from 15.2 million shipments in 2020 to 2.5 billion in 2030 [abi2021].
[…]
In terms of use cases, TinyML has been widely deployed for (i) word spotting, (ii) object recognition, (iii) object counting, and (iv) audio/voice detection [tinyml, abi2022]. More recently, its application has been extended to the detection of (i) forest fire, (ii) shape, and (iii) seizure [abi2022].
@@ 2.2 Scope
This work targets ultra-low-power ML applications (TinyML), commonly powered by Arm Cortex-M (Armv7-M or Armv8-M) and RISC-V MCUs (e.g., PULP [pulp]). These devices typically have a single core and feature Flash and static random access memories (SRAM). They execute instructions in place from physical memory and lack a memory management unit (MMU) for virtual memory support. Consequently, they are unable to run Unix-like operating systems.
@@
[…]
@@ 2.3 Adversary Model
Adversary Goal. The adversary’s goal is to compromise the decision integrity of a supervised model, disrupting severe accuracy drops. Previous research shows that the attacker can achieve this goal by interfering with the training or inference phase. However, research on training at the deep edge is still at an embryonary stage [costa2022], and this class of devices is still only used for inference purposes. The most relevant part of the training of the models deployed on deep-edge devices is still performed in cloud-like environments. Therefore, we limit the scope of our research to the inference phase. Attacks to craft adversarial examples can either be targeted or non-targeted. Targeted attacks misguide an ANN by making it classify an adversarial sample as a specific class. In contrast, in non-targeted attacks, the class returned by the model is arbitrary except for the original one. Our research encompasses only non-targeted attacks, as they enable a broader assessment of the robustness and generalization of a neural network.
Adversary Capability. Capability refers to the attacker’s capacity to control or manipulate the input data. At inference, the generic data processing pipeline of an ML system considers that the (i) input features are collected from sensors or data repositories and (ii) processed in the digital domain, (iii) used by the model to infer about the surrounding environment, (iv) which communicates the result to the appropriate actuator [papernot2018]. Considering this pipeline, adversarial examples can be generated either in the physical domain, by tampering with sensors, or in the digital domain. This work only considers attacks and defenses in the digital domain, which dominates the available literature.
@@ 2.4 Selected Datasets and Models
[…]
@@
We included a third dataset because evaluating attacks and defenses on only two datasets may fall short. For this purpose, we borrowed the work [lucan2022], which lists and reviews ML implementations for Arm Cortex-M MCUs. After analyzing the reproducibility of each listed work, we replicated the work [vita2020]. This work addressed the development of an ML system devoted to detecting coffee plant diseases. The dataset is public and contains 1747 images for training and testing distributed over four classes, each corresponding to a different disease. All QNNs included in our study are quantized post-training.
@@ 2.5 Selected Attacks and Defenses
We prioritize attacks/defenses with great impact and citation in the research community and whose implementation is open-sourced. Furthermore, since we target QNNs following the TF Lite Micro quantization policy, we prioritize attacks and defenses implemented in TensorFlow 2 (TF2) or Keras (the backend engine of TF2). TF2 provides straightforward mechanisms to operate with models in TF Lite format. We also prioritize attacks/defenses available in ART [art], Cleverhans [cleverhans], or Foolbox [foolbox1, foolbox2] frameworks, as these frameworks provide a straightforward mechanism to benchmark different attacks and defenses upon a common interface. Our evaluation suite includes six white-box, two gray-box, and two black-box attacks, described in Table II.
@@
Defenses. We consider six defense mechanisms (Table III) distributed across the two types of defense proposed by the research community. Our empirical evaluation includes four train-based defenses and the other two based on input pre-processing. Although there might be more recent defenses available in the literature than those we selected, PGD Adversarial Training, proposed in 2018 and included in our work, remains the state-of-the-art defense against adversarial attacks according to recent studies [autoattack, zhang2021_2, ferrari23]. Nevertheless, we still include a recent iteration of this defense, entitled Sinkhorn Adversarial Training (2021), which promises to reduce the ground distance between original images and their adversarial representations.
@@
Zeroth-Order Optimization (ZOO) is a gray-box version of C&WL2. ZOO uses an alternative formula to estimate the gradient, where $h$ is a small constant and $e_{i}$ is a standard basis vector with only the $i^{th}$ component set to 1. To be efficient, ZOO uses stochastic coordinate descent methods.
@@
Adversarial training consists in retraining an ANN by injecting adversarial examples into the training dataset. The first work on adversarial training [fgsm] proposed to generate adversarial examples using FGSM in each training iteration, mixing clean samples and adversarial examples. This work was later borrowed by [pgd], which proposed adversarial training with PGD.
@@ 2.7 Threats to Validity
The experiments were conducted by two independent researchers on two separate but equal platforms: STM NUCLEO-L552ZE-Q. Whenever possible, we used the implementation of attacks and defenses available on the ART framework [art], except for AutoAttack and Sinkhorn Adversarial Training. Regarding AutoAttack, the original implementation is significantly different from the implementation of ART. Sinkhorn Adversarial Training is not available on ART and is originally implemented in PyTorch. In these two cases, we borrowed the original implementations and adapted them to work with TF2. Despite all efforts, there may be slight variations compared to the original works. To encourage further reproducibility of our results, all artifacts are available open-source [^1] https://gitlab.com/ESRGv3/QNN-Adversarial-Robustness.git.
@@ 3.1 Direct Attacks
[…]
{mdframed}
[style=remarkstyle] Takeaway 1. The average point distance to the decision boundary increases as the precision of a neural network decreases, making int-8 models generally more robust to adversarial examples than int-16 and float-32 versions.
@@
[…]
bernhard2019 arrived at a similar conclusion. They observed that the quantization of activations could make the gradient explode or vanish. However, they did not observe the same behavior for weight-only quantized models, i.e., models with weights quantized to integer precision but whose activations remain in floating-point.
{mdframed}
[style=remarkstyle] Takeaway 2. Quantization increases the chance of gradient masking. Small perturbations may be erased by quantization and make the gradient vanish while bigger perturbations can make the gradient explode.
@@ 3.2 Transferability
A3. Transferability is sensitive to quantization-shift. As observed in Figure 1, for the CIFAR-10 and Coffee datasets, the lines representing the adversarial accuracy of float-32 and int-16 models overlap for the entire range of distortion. However, the same does not occur with the int-8 model. While for float-32 and int-16, the adversarial accuracy decreases smoothly as distortion increases, for the int-8 model, the adversarial accuracy decreases in buckets - it only decreases after a certain distortion threshold. We argue that this is a consequence of the quantization shift, the most widely-known effect of quantization [bernhard2019, gorsline2021, zhao2018, lin2019, duncan2020, song2021]. As quantization maps more than one value on the same quantization bucket, quantization can amplify or attenuate the effect of distortion. Quantization can boost adversarial accuracy by eliminating small perturbations; however, it can amplify bigger perturbations.
@@
{mdframed}
[style=remarkstyle] Takeaway 3. Adversarial example transferability is influenced by quantization-shift. Quantization can mitigate the impact of small perturbations while amplifying the effects of larger perturbations.
[…]
@@
{mdframed}
[style=remarkstyle] Takeaway 4. By mapping different ANN weights and activations into the same integer value, quantization can cause gradient misalignment: the gradient estimated for float-32 ANNs may not be the same for the equivalent int-16 or int-8 models and even have different directions.
@@ 4 Defenses Against Adversarial Examples
  • Gray-box: We only include the Square Attack as the results detailed in Table V suggest that ZOO does not perform well against the neural networks we are using.
@@ 4.1 Train-Based Defenses
{mdframed}
[style=remarkstyle] Takeaway 5. PGD and Sinkhorn Adversarial Training consistently increase the average point distance to the decision boundary, and this feature holds after quantization. Defensive Distillation and Ensemble Adversarial Training are not consistent in this topic.
@@
{mdframed}
[style=remarkstyle] Takeaway 6. None of the training-based defenses mitigate adversarial example transferability. We envision this is a consequence of none of them smoothing the quantization-shift and gradient misalignment, which are required to tackle the transferability.
@@
{mdframed}
[style=remarkstyle] Takeaway 7. Every time a defense is effective in a full-precision ANN, it is likewise effective in the corresponding QNNs. A defense may, however, be adequate to defend a low-precision QNN but insufficient to defend a full-precision ANN.
@@ 4.2 Pre-Processing Defenses
Pre-processing defenses are evaluated against the adversarial examples produced to attack the vanilla ANNs and QNNs before any defense. We do not craft new adversarial examples for the Defense+ANN or Defense+QNN system as pre-processing defenses are known to disrupt the well-known gradient-masking phenomenon, in which the gradient becomes useless to create an adversarial example, giving a false sense of security [bernhard2019, gorsline2021, lin2019]. The evaluation of Pixel Defend was restricted to CIFAR-10 due to hardware limitations: PixelCNN is a heavy network whose amount of neurons grows proportionally with the size of input images. Therefore, we limited the evaluation scope to the dataset with smaller images. The configuration of Feature Squeezing and Pixel Defend is detailed in Table XV (Appendix A), while the test accuracy of the ANNs/QNNs enhanced with these defenses is detailed in Table X. Tables XI and XII detail the results observed for Feature Squeezing and Pixel Defend, respectively.
@@
{mdframed}
[style=remarkstyle] Takeaway 8. Feature Squeezing can denoise a larger fraction of adversarial examples than Pixel Defend. In contrast to Pixel Defend, Feature Squeezing is portable to edge devices powered by Arm Cortex-M MCUs as long as they feature a DSP.
@@
{mdframed}
[style=remarkstyle] Takeaway 9. Feature Squeezing is effective against small perturbations, but the spatial smoothing filter starts to fall short as the amount of perturbed pixels increases.
@@
{mdframed}
[style=remarkstyle] Takeaway 10. Train-based defenses do not incur additional computation during inference and, therefore, are preferable over pre-processing defenses to be deployed in edge devices powered by Arm Cortex-M MCUs.